# Detecting Dark Matter Substructures on Small Scales with Fast Radio Bursts

### v1

@@ Abstract
The matter power spectrum is only weakly constrained on subgalactic scales, while physics beyond the Standard Model can leave unique imprints especially on sub-parsec scales. We propose measuring the arrival time difference of Fast Radio Bursts (FRBs) along two adjacent sightlines as a new probe to dark matter substructures on scales down to $\sim 1\,$AU. We discuss two observational scenarios in which it may be possible to place interesting constraints on such models through monitoring repeating FRB sources: 1) By sending radio receivers to space to form a baseline of tens of AU or more and measuring the temporal variation of the arrival time difference between receivers. 2) By measuring the temporal variation of the arrival time difference between two lensed images of one strongly lensed repeater. In both scenarios, obtaining interesting constraints requires correlating the voltage time series to measure the radio-signal arrival time to sub-nanosecond precision. We find that two radio dishes separated by 202020\,20AU may be sensitive to the enhancement of small-scale structures at $\sim 10^{-8}M_{\odot}$ masses in the QCD axion dark matter scenario or from an early epoch of matter-domination with a reheating temperature up to 60 MeV. Other dark matter models such as those composed of $\sim 10^{-13}M_{\odot}$ primordial black holes produced during inflation would also be probed by this method. We further show that a strong lensing situation of multiple images provides an equivalent $\sim 2000\,$AU baseline, which can be much more sensitive but with the uncertainty that intervening ISM decoherence may degrade the timing precision and that spatial variation in the FRB emission spot may result in confounding signals. We show that the lensing magnifications of Type Ia supernovea constrain a similar quantity to such FRB timing, with present limits being equivalent to ruling out the same parameter space that would be probed by a 0.140.140.14~{}0.14AU baseline.
@@ I Introduction
Dark Matter (DM) structures on comoving length scales shorter than tens of kiloparsecs all the way down to the solar-system scale have only been weakly constrained by observations [1, 2, 3, 4, 5, 6]. Enhanced formation of self-gravitating structures on such tiny scales than in the Cold Dark Matter (CDM) scenario may arise from special physical processes that take place in the early Universe. Examples include an early epoch of matter domination [7, 8], an axion arising from a Peccei-Quinn phase transition after inflation [9], a primordial kination era [10], inflationary production of vector DM [11] and density perturbations from magnetic fields [12]. Therefore, probing such minuscule DM structures would open up a window into the Universe before Big Bang Nucleosynthesis and provide opportunities to discover particle physics beyond the Standard Model. The only surefire way to detect such structures is via their gravitational effects.
@@
In this work, we propose the precision comparison of the arrival times of repeated Fast Radio Burst (FRB) signals along multiple sightlines from the same FRB source as a new direct probe for minuscule DM substructures. FRBs are radio transients with typically millisecond duration powered by mostly extragalactic sources. While the origin of FRBs is not yet thoroughly understood, thousands of them have been discovered [36], with over fifty currently known to repeat [37, often on tens of hour timescales], and their propagation effects provide a powerful tool to study cosmological physics such as the missing baryon problem, the circumgalactic media of galaxies, the cosmic reionization history [38, 39, 40, 41, 42], and, in this work, the dark matter substructures. This proposal uses two key aspects of FRBs: they originate from extremely compact sources, and relative timing of the radio signals can be performed to sub-nanosecond precision through coherent analysis of voltage time series [43].
@@ II Weak Lensing Effect by Dark Matter Substructures for Two-Dish Configuration
This simple estimate assumes the DM structure is smaller than $x_{0}$. The calculations we present later will properly account for the size of DM structures.
Our method can be compared to two other Galactic methods to measure small DM structures – pulsar timing arrays (PTAs) and astrometry in the proper motion of Galactic stars. One advantage of our method is the involvement of cosmological path lengths which leads to overall delays that are $\sim$ 100 times larger than in the Galaxy. PTAs are sensitive to residuals of $\sim 100\,$ns [55], and again we expect the Galactic contribution to our $\Delta t$ to be $\sim 1\%$ of the total delay so our hypothetical experiment should be more sensitive. To compare with astrometry, the 0.001 nanosecond timing accuracy that probes an equivalent structure as our mass corresponds to $\mu$arcsecond deflections of stars using that $\Delta t\sim d\,\Delta\theta^{2}/2$ and a galactic distance of $d=$ 1 kpc. Future surveys like the Square Kilometre Array (SKA) can reach an angular resolution of 10 $\mu$arcseconds or even $\mu$arcsecond to the brightest stars with 10 years of operation [56]. Another advantage of probing minuscule DM structures with cosmological sources such as FRBs is that those DM structures are less disrupted in the low-density intracluster [16]) or intergalactic space, but more so in the Galaxy due to tidal stripping and stellar disruptions [57, 58]. The microlensing of magnified stars proposal also uses extragalactic sources [16] but it is challenging to access its sensitivity and currently there is no estimation.
@@ II.1 Arrival time difference introduced by dark matter substructures
A larger transverse separation $x_{0}$ will lead to a more detectable effect (the time delay signal is proportional to $x_{0}^{2}$). The overall contribution of this higher-order term (the quadrupole term or the circularly symmetric term) in Eq. 11 where the delay scales quadraticly in $|x_{\perp}|$ is suppressed due to the small dish separation. However, this suppression can be alleviated by looking at the temporal change on the arrival time difference if we detect repeating FRBs, as we will see in the next subsection. In other words, we can use the measurement of repeating FRBs to break the degeneracy on the leading order with source angle. Additionally, the movement of dark matter substructures can effectively increase the baseline length and the variance of the arrival time difference can be enhanced by a factor of $(x_{0}/[v\,\delta t])^{2}$, where $\delta t$ is the duration over which the repeating FRB is observed and $v$ is the velocity of dark matter substructures. Even though observing the quartic contribution in Eq. 10 is less sensitive except for extremely large baselines, it has the advantage that it does not require the FRB to repeat and can be used as a cross check of the method we discuss next.
@@ II.2 Time variance of weak lensing effect with repeating FRBs
[…]
@@
If potential fluctuation modes of short wavelengths are not a dominant contribution, i.e. if $k_{\perp}\sigma_{v}\,\delta t\ll 1$, the time delay difference $\Delta t$ between the two sightlines as a function of time $t$ can be expanded as a Taylor series $\Delta t=a_{0}+a_{1}\,t+a_{2}\,t^{2}+...$. The expansion coefficients, encoding information about the DM substructures, can be measured accurately with sufficiently many well-measured FRB repetitions. The leading term (constant term $a_{0}$), independent of $\delta t$, is described by Eq. 9 when expanding the Bessel function, which is degenerate with a linear shift on the source angular location. The next term is related to the variance of the time derivative on the arrival time difference, which is expressed as
@@
In Fig. 2, we plot the sensitivity on the matter power spectrum for the 2-Dish configuration, from measuring ${\rm d}\Delta t/{\rm d}t$ and ${\rm d}^{2}\Delta t/{\rm d}t^{2}$, where $\Delta t$ is the arrival time difference along two sightlines. The 2-Dish system refers to radio receivers in space that create separate sightlines for the same FRB source and separated by $x_{0}=20$AU (left) and $x_{0}=0.14$AU (right). The sensitivity curves are calculated for the following fiducial parameter values: FRB timing precision $\delta t_{m}=0.1$ ns, comoving distance to the source of $D_{\rm s}=3$ Gpc (or $z\approx 1$), characteristic dispersion of peculiar velocity of the DM structures $\sigma_{v}=500$ km/s, time span of observing repeating FRB sources $\delta t=10$ yr, and the number of $\delta t_{m}$-timed radio bursts is $N=100$. The matter power spectrum expected from the standard $\Lambda$CDM cosmology is also plotted in Fig. 2, though we must emphasize that the high-$k$ regime ($k>10\,\rm Mpc^{-1}$) is poorly constrained. The black dashed curve in Fig. 2 is the linear matter power spectrum in $\Lambda$CDM cosmology, while the blue dashed curve is the nonlinear matter power spectrum obtained from the HaloFit model [64, 65]. Our observational probe constrains the nonlinear power spectrum. The blue region is excluded by current observations from CMB [61, 66, 67, 62], and Lyman-alpha forest [63] (The blue triangular region is currently excluded by limits on CMB spectral distortions; ). The gray region is the sensitivity forecast for the 2-Dish system we study in this work. Technically this system requires four dishes to constrain their positions via GPS techniques, although the other dishes are not necessarily as large as the “2-Dish” we are relying on to detect FRBs. The sensitivities of the 2-Dish system to ${\rm d}\Delta t/{\rm d}t$ and ${\rm d}^{2}\Delta t/{\rm d}t^{2}$ are respectively calculated using Eq. 17 and Eq. 21.
@@
Large scale structures in the $\Lambda$CDM cosmology that have 10s or even hundreds of wavelengths still affect the first time derivative of arrival time difference from the gradient field of their gravitational potential (as seen by the 2-Dish linear running somewhat parallel to the matter power spectrum in Fig. 2). This means that baselines much longer than $x_{0}=0.14$ will detect the $\Lambda$CDM contribution and not be able to probe the potential high-$k$ enhancements we are interested in. The way to avoid this large-scale sensitivity is to measure the second time derivative of the arrival time difference, as it will only be sensitive to the matter overdensity modes at higher $k$ with the $\Lambda$CDM contribution efficiently filtered out. The sensitivity to the second time derivative is shown by the orange curves in Fig. 2.
[…]
@@ III.1 Weak Lensing Effects In Strong Lensing Systems
The above expression can be compared to the sensitivity in the 2-Dish system discussed in Eq. 17. Therefore, the sensitivity of the 2-Image configuration is comparable to a 2-Dish configuration with a 2000 AU baseline. Note that the second derivative term in the 2-Image scenario is equivalent to the first derivative term in the 2-Dish system. Likewise, when we expand the 2-Image signal to the third derivative (cubic term), it will correspond to the quadratic term discussed in Eq. 21. Due to the great sensitivity, the signal might be contaminated by $\Lambda$CDM large scale structures, which can be mitigated by measuring higher-order time derivatives of the arrival time difference.
@@ III.2 Decoherence and Scattering Effects Between Sightlines
Realizing the method we propose requires that the FRB radio signals that travel along different sightlines have phase coherent voltage time series to allow cross-correlation. We now examine the plausibility of this assumption for both the 2-Dish configuration and the 2-Image scenario. The most outstanding issue of phase decoherence likely comes from the interstellar medium (ISM) of the Milky Way. ISM can cause pulse delay and broadening along sightlines to individual radio receivers in space. This is known as the scattering effect, which is induced by density inhomogeneities in the interstellar plasma. Observations of pulsars and FRBs away from the Galactic disk suggest a scattering time $\delta t_{\rm sc}\simeq 30\,{\rm ns}\,(\nu/\rm GHz)^{-4}\,(\sin b)^{-2.5}$, where $b$ is the Galactic latitude of the sightline [e.g. 48]. Pulse delay and broadening are typically on this timescale. If the scattering-induced time delay is significant ($\delta t_{\rm sc}>1/\nu$), radio waves travel along multiple geometric paths with significant phase differences, resulting in decoherence in the voltage time series. To have $\delta t_{\rm sc}<\nu^{-1}$, observations need to be carried out at $\nu\gtrsim 3$ GHz to suppress ISM decoherence. For our fiducial 0.10.10.1\,0.1ns timing accuracy, we would require $\delta t_{\rm sc}<0.1\,{\rm ns}$; otherwise the timing error worsens by scattering. See Ref. [48] for more extensive discussions on the scattering effects on FRB observations in the context of space interferometry on the Solar System scale. Finally, scattering in the lens itself could be important since the light paths may pass close to the denser lens center and its host galaxy’s ISM.

### v2

@@ Abstract
The matter power spectrum is only weakly constrained on subgalactic scales, while physics beyond the Standard Model can leave unique imprints especially on sub-parsec scales. We propose measuring the arrival time difference of Fast Radio Bursts (FRBs) along two adjacent sightlines as a new probe to dark matter substructures on scales down to $\sim 1\,$AU. We discuss two observational scenarios in which it may be possible to place interesting constraints on such models through monitoring repeating FRB sources: 1) By sending radio receivers to space to form a baseline of tens of AU or more and measuring the temporal variation of the arrival time difference between receivers. 2) By measuring the temporal variation of the arrival time difference between two lensed images of one strongly lensed repeater. In both scenarios, obtaining interesting constraints requires correlating the voltage time series to measure the radio-signal arrival time to sub-nanosecond precision. We find that two radio dishes separated by 202020\,20AU may be sensitive to the enhancement of small-scale structures at $\sim 10^{-8}M_{\odot}$ masses in the QCD axion dark matter scenario or from an early epoch of matter-domination with a reheating temperature up to 60 MeV. Other dark matter models such as those composed of $\sim 10^{-13}M_{\odot}$ primordial black holes produced during inflation would also be probed by this method. We further show that a strong lensing situation of multiple images provides an equivalent $\sim 2000\,{\rm AU}\,(\sigma_{v}/10^{3}\,{\rm km}{\rm s}^{-1})\,(\delta t/10\,{\rm yr})$ baseline, for a typical velocity of DM substructures $\sigma_{v}$ and an observational time span $\delta$. This is much more sensitive but with the uncertainty that intervening ISM decoherence may degrade the timing precision and that spatial variation in the FRB emission spot may result in confounding signals. We show that the lensing magnifications of Type Ia supernovea constrain a similar quantity to such FRB timing, with present limits being equivalent to ruling out the same parameter space that would be probed by a 0.140.140.14~{}0.14AU baseline.
@@ I Introduction
Dark Matter (DM) structures on comoving length scales shorter than tens of kiloparsecs all the way down to the solar-system scale have only been weakly constrained by observations [1, 2, 3, 4, 5, 6]. Enhanced formation of self-gravitating structures on these tiny scales than in the Cold Dark Matter (CDM) scenario may arise from special physical processes that take place in the early Universe. Examples include an early epoch of matter domination [7, 8], an axion arising from a Peccei-Quinn phase transition after inflation [9], a primordial kination era [10], inflationary production of vector DM [11] and density perturbations from magnetic fields [12, 13]. Therefore, probing such minuscule DM structures would open up a window into the Universe before Big Bang Nucleosynthesis and provide opportunities to discover particle physics beyond the Standard Model. The only surefire way to detect such structures is via their gravitational effects.
@@
In this work, we propose the precision comparison of the arrival times of repeated Fast Radio Burst (FRB) signals along multiple sightlines from the same FRB source as a new direct probe for minuscule DM substructures. FRBs are radio transients with typically millisecond duration powered by mostly extragalactic sources. While the origin of FRBs is not yet thoroughly understood, thousands of them have been discovered [37], with over fifty currently known to repeat [38, often on tens of hour timescales], and their propagation effects provide a powerful tool to study cosmological physics such as the missing baryon problem, the circumgalactic media of galaxies, the cosmic reionization history [39, 40, 41, 42, 43], and, in this work, the dark matter substructures. This proposal uses two key aspects of FRBs: they originate from extremely compact sources, and relative timing of the radio signals can be performed to sub-nanosecond precision through coherent analysis of voltage time series [44].
@@ II Weak Lensing Effect by Dark Matter Substructures for Two-Dish Configuration
This simple estimate assumes the DM structure is smaller than $x_{0}$. The calculations we present later will properly account for the size of DM structures. Our method can be compared to two other Galactic methods to measure small DM structures – pulsar timing arrays (PTAs) and astrometry in the proper motion of Galactic stars. One advantage of our method is the involvement of cosmological path lengths which leads to overall delays that are $\sim$ 100 times larger than in the Galaxy. PTAs are sensitive to residuals of $\sim 100\,$ns [56], and again we expect the Galactic contribution to our $\Delta t$ to be $\sim 1\%$ of the total delay so our hypothetical experiment should be more sensitive. For PTAs, astrophysical systematics such as the spin noise of pulsars and the time variation of electron density (or dispersion measure) induced by the solar wind will have a great impact on the timing precision [57, 58]. However, FRB timing proposed in this work is intrinsically different from pulsar timing because pulsar timing relies on pulsars being a reference clock while FRB timing uses the difference in arrival times between detectors at the end of different sightlines. By correlating the electric field at two different receivers, the dispersion measure difference, which only enters the correlation as a phase, can be fitted away with a broad frequency band. Therefore, those astrophysical systematics that arise from pulsars or solar wind will not be a concern in our proposal of FRB timing.
To compare with astrometry, the deflection angle is $\Delta\theta\sim\delta t/b$ where $b\sim 100$ AU is the typical impact parameter and $\delta t$ is the Shapiro time delay induced by dark matter substructures. Therefore, a micro-arcsecond precision on the deflection is equivalent to a timing precision of $100$ ns. In addition, astrometry uses dark matter substructures in the galaxy so the number of objects will be smaller than the estimate for cosmological sources. Future surveys like the Square Kilometre Array (SKA) can reach an angular resolution of 10 $\mu$arcseconds or even $\mu$arcsecond to the brightest stars with 10 years of operation [59]. Another advantage of probing minuscule DM structures with cosmological sources such as FRBs is that those DM structures are less disrupted in the low-density intracluster [17] or intergalactic space, but more so in the Galaxy due to tidal stripping and stellar disruptions [60, 61]. The microlensing of magnified stars proposal also uses extragalactic sources [17] but it is challenging to access its sensitivity and currently there is no estimation.
@@ II.1 Arrival time difference introduced by dark matter substructures
A larger transverse separation $x_{0}$ will lead to a more detectable effect (the time delay signal is proportional to $x_{0}^{2}$). The overall contribution of this higher-order term (the quadrupole term or the circularly symmetric term) in Eq. 11 where the delay scales quadraticly in $|x_{\perp}|$ is suppressed due to the small dish separation. However, this suppression can be alleviated by looking at the temporal change on the arrival time difference if we detect repeating FRBs, as we will see in the next subsection. In other words, we can use the measurement of repeating FRBs to break the degeneracy on the leading order with source angle. [^2] This does additionally require an absolute measurement of the angle to the FRB, which necessitates knowing the overall rotational orientation of the array and not just the relative positions of the receivers. This requires using other FRBs to set the reference frame or, possibly, precise two-way ranging to measure the distances between the array elements and terrestrial receivers whose position is extremely well modeled. Additionally, the movement of dark matter substructures can effectively increase the baseline length and the variance of the arrival time difference can be enhanced by a factor of $(x_{0}/[v\,\delta t])^{2}$, where $\delta t$ is the duration over which the repeating FRB is observed and $v$ is the velocity of dark matter substructures. Even though observing the quartic contribution in Eq. 10 is less sensitive except for extremely large baselines, it has the advantage that it does not require the FRB to repeat and can be used as a cross check of the method we discuss next.
@@ II.2 Time variance of weak lensing effect with repeating FRBs
In reality, the DM substructure velocity $\bm{v}_{\perp}$ involved in the above expression varies along the sightline. The peculiar motion of the large-scale structure is coherent over distances of tens of megaparsecs, while the FRB source is at a cosmological distance. The sightline therefore samples different coherent patches of the large-scale velocity field. This justifies taking the statistical average of the above expression with respect to the distribution of $\bm{v}_{\perp}$. Accounting for large-scale linear matter overdensity modes, the velocity is drawn from a Gaussian distribution with a standard deviation of $\sigma_{v}\approx 600\sqrt{2/3}$ km s-1 at low redshifts, where $\sqrt{2/3}$ accounts for only the transverse component of the velocity vector  [^3] We neglect the contribution from local collapsed structures here, which would increase somewhat the dispersion.. By integrating $\sigma_{t}^{2}(v_{\perp})$ over the Gaussian distribution of $v_{\perp}$, we derive:
@@
If potential fluctuation modes of short wavelengths are not a dominant contribution, i.e. if $k_{\perp}\sigma_{v}\,\delta t\ll 1$, the time delay difference $\Delta t$ between the two sightlines as a function of time $t$ can be expanded as a Taylor series $\Delta t=a_{0}+a_{1}\,t+a_{2}\,t^{2}+...$. The expansion coefficients, encoding information about the DM substructures, can be measured accurately with sufficiently many well-measured FRB repetitions. The leading term (constant term $a_{0}$), independent of $\delta t$, is described by Eq. 9 when expanding the Bessel function, which is degenerate with the source angular location. The next term is related to the variance of the time derivative on the arrival time difference, which is expressed as
@@
In Fig. 2, we plot the sensitivity on the matter power spectrum for the 2-Dish configuration, from measuring ${\rm d}\Delta t/{\rm d}t$ and ${\rm d}^{2}\Delta t/{\rm d}t^{2}$, where $\Delta t$ is the arrival time difference along two sightlines. The 2-Dish system refers to radio receivers in space that create separate sightlines for the same FRB source and separated by $x_{0}=20$AU (left) and $x_{0}=0.14$AU (right). The sensitivity curves are calculated for the following fiducial parameter values: FRB timing precision $\delta t_{m}=0.1$ ns, comoving distance to the source of $D_{\rm s}=3$ Gpc (or $z\approx 1$), characteristic dispersion of peculiar velocity of the DM structures $\sigma_{v}=500$ km/s, time span of observing repeating FRB sources $\delta t=10$ yr, and the number of $\delta t_{m}$-timed radio bursts is $N=100$. The matter power spectrum expected from the standard $\Lambda$CDM cosmology is also plotted in Fig. 2, though we must emphasize that the high-$k$ regime ($k>10\,\rm Mpc^{-1}$) is poorly constrained. The black dashed curve in Fig. 2 is the linear matter power spectrum in $\Lambda$CDM cosmology, while the blue dashed curve is the nonlinear matter power spectrum obtained from the HaloFit model [67, 68]. Note that the nonlinear power spectrum has been extrapolated to an extremely high wavenumber regime, well beyond the wavenumbers the HaloFit model has not been calibrated to simulations. Our observational probe constrains the nonlinear power spectrum. The blue region is excluded by current observations from CMB [64, 69, 70, 65], and Lyman-alpha forest [66] (The blue triangular region is currently excluded by limits on CMB spectral distortions; ). The gray region is the sensitivity forecast for the 2-Dish system we study in this work. Technically this system requires four dishes to constrain their positions via GPS techniques, although the other dishes are not necessarily as large as the “2-Dish” we are relying on to detect FRBs. The sensitivities of the 2-Dish system to ${\rm d}\Delta t/{\rm d}t$ and ${\rm d}^{2}\Delta t/{\rm d}t^{2}$ are respectively calculated using Eq. 17 and Eq. 21.
@@
Large scale structures in the $\Lambda$CDM cosmology that have 10s or even hundreds of wavelengths still affect the first time derivative of arrival time difference from the gradient field of their gravitational potential (as seen by the 2-Dish linear running somewhat parallel to the matter power spectrum in Fig. 2). This means that baselines much longer than $x_{0}=0.14\,$AU will detect the $\Lambda$CDM contribution and not be able to probe the potential high-$k$ enhancements we are interested in. The way to avoid this large-scale sensitivity is to measure the second time derivative of the arrival time difference, as it will only be sensitive to the matter overdensity modes at higher $k$ with the $\Lambda$CDM contribution efficiently filtered out. The sensitivity to the second time derivative is shown by the orange curves in Fig. 2.
The expansion of the time delay difference in terms of the first derivatives, the second derivative, or even higher-order derivatives, is only valid when the dominant modes satisfy $k\,\sigma_{v}\,\delta t\ll 1$. When $k$ is larger than $1/(\sigma_{v}\,\delta t)$, the temporal power spectrum of the arrival time difference is the appropriate observable that preserves the full information. This change over in constraining method is reflected by the kinks in the sensitivity curves, where the higher $k$ beyond the kink uses the temporal power spectrum to constrain the matter power spectrum (which will be discussed in the next subsection). Practically, the highest wavenumber that such FRB timing may probe is given by repeating rate of FRBs, which for some of the most prolific repeaters is $\sim 1/(10)~{}$hr-1 [63], divided by the velocity of the dark matter substructure, $\sigma_{v}$, which yields $k\sim 10^{12}\,\rm Mpc^{-1}$.
@@ III.1 Weak Lensing Effects In Strong Lensing Systems
The above expression can be compared to the sensitivity coming from the first time derivative of the arrival time variation for the 2-Dish system, as discussed in Eq. 17. The sensitivity of the 2-Image configuration is equivalent to a 2-Dish configuration with a baseline of $2000\,{\rm AU}\,(\delta t/10\,{\rm yr})\,(\sigma_{v}/10^{3}\,{\rm km}{\rm s}^{-1})$. Note that the second time derivative term in the 2-Image scenario has the same $k$ scaling as the first derivative term in the 2-Dish system. Likewise, when we expand the 2-Image signal to the third time derivative (cubic term), it has the same $k$ scaling as the second time derivative term discussed in Eq. 21. The difference in the $k$ scaling, for $k\,x_{0}<1$, results from the difference that the two slightlines to the two receiver dishes probe similar potential fluctuations while in the strong lensing scenario potential fluctuations near the two widely-separated light paths are uncorrelated. Lower time derivatives of the arrival time difference are more contaminated by long wavelength modes in the $\Lambda$CDM scenario. This limitation can be circumvented by measuring higher-order time derivatives of the arrival time difference.
@@ III.2 Decoherence and Scattering Effects Between Sightlines
Realizing the method we propose requires that the FRB radio signals that travel along different sightlines have phase coherent voltage time series to allow cross-correlation. We now examine the plausibility of this assumption for both the 2-Dish configuration and the 2-Image scenario. The most outstanding issue of phase decoherence likely comes from the interstellar medium (ISM) of the Milky Way. ISM can cause pulse delay and broadening along sightlines to individual radio receivers in space. This is known as the scattering effect, which is induced by density inhomogeneities in the interstellar plasma. Observations of pulsars and FRBs away from the Galactic disk suggest a scattering time $\delta t_{\rm sc}\simeq 30\,{\rm ns}\,(\nu/\rm GHz)^{-4}\,(\sin b)^{-2.5}$, where $b$ is the Galactic latitude of the sightline [e.g. 49]. Pulse delay and broadening are typically on this timescale. If the scattering-induced time delay is significant ($\delta t_{\rm sc}>1/\nu$), radio waves travel along multiple geometric paths with significant phase differences, resulting in decoherence in the voltage time series. To have $\delta t_{\rm sc}<\nu^{-1}$, observations need to be carried out at $\nu\gtrsim 3$ GHz to suppress ISM decoherence. For our fiducial 0.10.10.1\,0.1ns timing accuracy, we would require $\delta t_{\rm sc}<0.1\,{\rm ns}$; otherwise the timing error worsens by scattering. See Ref. [49] for more extensive discussions on the scattering effects on FRB observations in the context of space interferometry on the Solar System scale. Finally, scattering in the lens itself could be important since the light paths may pass close to the denser lens center and its host galaxy’s ISM.